Consider a car moving on a straight road with a speed of 100 metre per second the distance at which the car can be stop is UK is equals to 0.5

Grade:12

1 Answers

Shashank
6 years ago
Given :
u = 100 m/s
v = 0 m/s
uk = 0.5
 
Solution :
Fk = uk * N
ma = uk * mg
ma = uk * mg
a = 0.5 * 10 m/s2
a = 5 m/s2
 
v = u + at
0 = 100 + (-5)t
t = 100/5 s
t = 20 s
 
s = ut + ½ at2
s = (100)(20) + ½ (-5)(20)2 m
s = 2000 - ½ (2000) m
s = 2000 – 1000 m
s = 1000 m